Melissa is 54 years old and her mother, Katie is 80. How many years ago was her mother, three times the age of her daughter, Melissa?
Answer by mananth(16949)

Melissa = 54
Katie = 80
let her mother be 3 times x years ago
(80-x)=3(54-x)
80-x=162-3x
2x=162-80
2x=82
x=41
Check
Katie = 80-41=39
Melissa=54-41=13